OpenStudy (anonymous):

Jacob and Roman are standing next to each other. Jacob, who is one hundred eighty centimeters tall, has a shadow that is forty-five centimeters long. If Roman is one hundred sixty-eight centimeters tall, approximately how long is his shadow?

OpenStudy (anonymous):

A)28 B)6 C)18 D)42

OpenStudy (welshfella):

similar triangles 45 / 180 = x / 168 find x

OpenStudy (welshfella):

180x = 45 * 168 x = (45 * 168) / 180

OpenStudy (anonymous):

@welshfella thank you :D

OpenStudy (perl):

test

OpenStudy (anonymous):

What do u mean?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

perl what do u mean by "test"